OpenStudy (anonymous):

Mia deposits either $5 or $3.50 into her savings account. She made 14 deposits equaling a total of $59.50. How many times did she deposit $5 and how many times did she deposit $3.50? Mia made deposits of $5 and deposits of $3.50.

OpenStudy (sleepyjess):

5x + 3.5y = 59.5 x + y = 14

OpenStudy (sleepyjess):

That is our system of equations

OpenStudy (anonymous):

@CanadianSteve99 @Tootles143

OpenStudy (anonymous):

Oh okay

OpenStudy (agl202):

the let call x number of deposits for $5 and y ......................................... for $3.5 total deposits x+y=14 5x number of deposit for $5 3.5y ....................... for $3.50 then the second equation is 5x+3.5y=59.50 substitute x with 14-y in the second equa to find x and then find y later x+y=14 5x+3.5y=59.50

OpenStudy (sleepyjess):

If you go to www.desmos.com and put in those 2 equations, see where they intersect

OpenStudy (anonymous):

so it is 14 ND 59.50??

OpenStudy (anonymous):

@sleepyjess Yeah i use desmos a lot but this question was a little confusing

OpenStudy (sleepyjess):

Yes, because they give us that there were 14 deposits and the deposits totaled 59.5

OpenStudy (anonymous):

do put the zero with 59.50?

OpenStudy (sleepyjess):

It doesn't matter

OpenStudy (anonymous):

A polygon is shown: A polygon MNOPQR is shown. The top vertex on the left is labeled M, and rest of the vertices are labeled clockwise starting from the top left vertex labeled, M. The side MN is parallel to side QR. The side MR is parallel to side PQ. The side MN is labeled as 5 units. The side QR is labeled as 7 units. The side MR is labeled as 3 units, and the side NO is labeled as 2 units. The area of polygon MNOPQR = Area of a rectangle that is 15 square units + Area of a rectangle that is ___ square units. (Input whole numbers only, such as 8.)

OpenStudy (sleepyjess):

..........

OpenStudy (sleepyjess):

Please don't spam others' posts with your question.

OpenStudy (anonymous):

You have aby clue @CanadianSteve99

OpenStudy (canadiansteve99):

No idea. sorry Typno, but I'll see if I can work it out